1/8 : x = 1/2 : 4/5
Answer by abdulaleem(11)

1/8 : x = 1/2 : 4/5
1/8 x 4/5 = x x 1/2
1/10 = x/2
1/10 x 2 = x
1/5 = x
so x = 1/5 answer